OpenStudy (anonymous):

Evaluate the integral (e^x)/((e^2x)+3) from 0 to infinity

OpenStudy (anonymous):

u = e^x du/dx = e^x du = e^xdx the integral becomes : du/(u^2 + 3) with boundaries 1 to infinity since e^0 = 1. the solution for this integral will be (1/(3^0.5)) * arctan(u/(3^0.5)) if you plug the boundaries you will get: (1/(3^0.5)) * (90 - 30) = 60/(3^0.5) hopefully i'm right lol

OpenStudy (anonymous):

.6046

OpenStudy (anonymous):

pi/(3 sqrt(3))=.6046

OpenStudy (anonymous):

integral (e^x)/((e^2x)+3) from 0 to infinity u = e^x +3 du = e^x dx = e^xdx the integral becomes =integral du/u from o to infinity = lnu] from 0 to infinity = ln(e^x +3)]from 0 to infinity

OpenStudy (anonymous):

so what is that value ? is it converges or divereges ?

OpenStudy (anonymous):

if im right its converges

OpenStudy (anonymous):

=0.6046 converge

OpenStudy (anonymous):

yep in radians

OpenStudy (anonymous):

so what is that value ? is it converges or divereges ?

OpenStudy (anonymous):

i think the u subst is wrong, becuz is e^2x